# Full exceptional collections on Fano threefolds and groups generated by spherical twists on K3 surfaces

### v3

@@ 7.1. $Y\subset{\mathbb{P}}^{3}$ (degree 4)
$-\otimes{\mathcal{O}}_{Y}(1)\in\operatorname{Aut}{\mathcal{D}}^{b}(Y).$
@@
In particular, $[\Gamma_{0}^{+}(2):G]=4$ and it is easy to see that $G$ is the free product of $4$ copies of ${\mathbb{Z}}/2{\mathbb{Z}}$, generated by $s_{0},ts_{0}t^{-1},t^{2}s_{0}t^{-2}$ and $t^{3}s_{0}t^{-3}$. These elements correspond to the spherical twists with respect to exactly the restrictions of the vector bundles of the exceptional collection ${\mathcal{F}}$. Figure 1 depicts a fundamental domain for $G$ which contains the corresponding $4$ fixed points of these generators.
@@ 7.2. $Y\subset Q_{3}$ (degree 6)
The group $\Gamma_{0}^{+}(3)\cong\operatorname{SO}^{+}(q,{\mathbb{Z}})$ has one elliptic point of order $2$, one elliptic point of order $6$ and one cusp ([16], Table 4). $X_{0}^{+}(3)$ has genus 0. Again, let $s_{0}\in\Gamma_{0}^{+}(3)$ be the transformation corresponding to $T_{{\mathcal{O}}_{Y}}\in\operatorname{Aut}{\mathcal{D}}^{b}(Y)$ and $t\in\Gamma_{0}^{+}(3)$ corresponding to $-\otimes{\mathcal{O}}_{Y}(1)\in\operatorname{Aut}{\mathcal{D}}^{b}(Y)$. Then, similarly to the previous example, by Proposition 3.13 we have $(ts_{0})^{3}=s$, where $s$ denotes the transformation corresponding to $T_{{\mathcal{S}}^{*}|_{Y}}$. Hence $ts_{0}$ generates an elliptic subgroup of order $6$. Same as in the previous example, up to an even shift, the autoequivalence ${\mathcal{O}}_{Y}(1)\otimes T_{{\mathcal{O}}_{Y}}(-)\in\operatorname{Aut}{\mathcal{D}}^{b}(Y)$ has order $6$ as well. The subgroup $G$ generated by reflections has index $3$ in $\Gamma_{0}^{+}(3)$:
@@ 7.3. $Y\subset V_{5}$ (degree 10)
$1\to G\to\Gamma_{0}^{+}(5)\cong({\mathbb{Z}}/2{\mathbb{Z}})^{\ast 3}\xrightarrow{s_{0}\mapsto 0,s_{1}\mapsto 0,t^{2}\mapsto 0}{\mathbb{Z}}/2{\mathbb{Z}}\to 1$
@@ 7.4. $Y\subset V_{22}$ (degree 22)
The group $\Gamma_{0}^{+}(11)\cong\operatorname{SO}^{+}(q,{\mathbb{Z}})$ has $4$ elliptic points of order $2$ and one cusp ([16], Table 4). $X_{0}^{+}(11)$ has genus 00.
@@ 7.5. Double cover of ${\mathbb{P}}^{2}$ (degree 2)
One observes that $[\Gamma_{0}^{+}(1):G]=3$ and $G\cong{\mathbb{Z}}/2{\mathbb{Z}}\ast{\mathbb{Z}}/2{\mathbb{Z}}\ast{\mathbb{Z}}/2{\mathbb{Z}}$, generated by $s_{0}$, $ts_{0}t^{-1}$ and $t^{2}s_{0}t^{-2}$. These elements correspond to spherical twists with respect to the pullbacks of the bundles forming the standard exceptional collection on ${\mathbb{P}}^{2}$. Figure 5 shows a fundamental domain for $G$ containing the corresponding $3$ fixed points.
@@ 7.6. An example of an infinitely generated group of spherical twists (degree 8)
The group ${\mathcal{G}}$ can be finitely generated even if the K3 surface in question is an anticanonical section of a Fano threefold without a full exceptional collection. For instance, let $X$ be a section of the Plücker embedding of $\operatorname{Gr}(2,6)$ by a subspace of codimension $5$. Then $X$ is a Fano threefold of Picard rank $1$, index $1$ and degree $14$ and a smooth anticanonical section $Y\subset X$ is a K3 surface of degree $14$. For such $Y$ the group ${\mathcal{G}}$ is finitely generated by Corollary 5.11, because $\Gamma_{0}^{+}(7)$ has genus 00, $2$ elliptic points of order $2$, which are both root points, $1$ elliptic point of order $3$ and $1$ cusp ([16, Table 4]). On the other hand, $X$ has no full exceptional collection since its Hodge diamond is not diagonal ($h^{2,1}=5$, see [26, §12.2]).
@@ 8. Exceptional collections on Fano threefolds
Let $X$ be a Fano threefold with Picard rank $1$ and very ample anticanonical class. Let $E,E^{\prime}$ be exceptional vector bundles on $X$ with the same class in $K_{0}(X)$. Assume that $\operatorname{Ext}^{1}(E,E\otimes\omega_{X}^{-1})=0$. Then $E\cong E^{\prime}$.
@@
By Lemma 3.11 we have $i^{*}E\cong i^{*}F[n]$. By shifting $E$ we may assume without loss of generality assume that $n=0$. Having done this it remains to show that $E$ has one nonzero cohomology sheaf in degree 00. Consider the distinguished triangle
@@
    Since $ad-bc$ is a unit and $a,d\in{\mathbb{Z}}_{p}$ we must have $v_{p}(b)+v_{p}(c)\geq 0$. Assume first $v_{p}(b)+v_{p}(c)>0$.
@@
    We now get that $k=2m-1$ (in particular, $k\neq 0$ and hence $p\mid\delta$) and $v_{p}(c)=m$, $v_{p}(b)=1-m$. So
@@ Appendix B Understanding $\operatorname{Aut}^{+}\operatorname{H}^{\ast}(Y)$
[22, Corollary 15.2.12] A K3 surface $Y$ of Picard rank $1$ has a non-trivial automorphism group if and only if it has degree $2$. In the latter case there is a (unique) nontrivial automorphism $\iota\in\operatorname{Aut}(Y)$, which is the covering involution. The involution $\iota$ acts as the identity on ${\mathcal{N}}(Y)$ and as $-1$ on the transcendental lattice $T(Y)={\mathcal{N}}(Y)^{\perp}\subset H^{2}(Y,{\mathbb{Z}})$.
@@
Let $Y$ be a K3 surface with Picard rank 1 and $d\neq 2$. Then there is no element $\iota$ in $\operatorname{Aut}^{+}H^{\ast}(Y,{\mathbb{Z}})$ which acts as the identity on ${\mathcal{N}}(Y)$ and as $-1$ on $T(Y)$.
@@
Assume $\iota$ exists. It follows from the definition that $\iota$ preserves the grading on $H^{\ast}(Y,{\mathbb{Z}})$. In particular it acts on $H^{2}(Y,{\mathbb{Z}})$ such that $\iota(H)=H$ and $\iota|_{T(Y)}=-1$. It follows that $-\iota$ is a reflection on $H^{2}(Y,{\mathbb{Z}})$ corresponding to $H$ and so it is given by the standard formula
@@
Let $Y$ be a K3 surface of odd Picard rank. The only Hodge isometries of the transcendental lattice $T(Y)$ are ±$\pm{\operatorname{id}}$.
@@
An element of $\operatorname{Aut}^{+}H^{*}(Y,{\mathbb{Z}})$ acts on the transcendental lattice $T(Y)$ by ±$\pm\operatorname{id}$, according to Proposition B.4. Since ±$\pm{\operatorname{id}}$ does not change the orientation of the positive 2-planes in $T(Y)_{\mathbb{R}}$, the map (B.1) is indeed well-defined.
If $\phi$ is a non-trivial element in the kernel of (B.1) then it is the identity on ${\mathcal{N}}(Y)$, and it acts as $-1$ on $T(Y)$. Invoking Proposition B.1 and Lemma B.3 finishes the proof. ∎
@@
which acts as ±$\pm 1$ on $T(Y)$. Note that by Lemma B.5 such an extension is unique for $d>2$ and for $d=2$ there is a ${\mathbb{Z}}/2{\mathbb{Z}}$ ambiguity which is inconsequential for what follows.
@@
    The kernel of $\operatorname{Aut}^{+}{\mathcal{N}}(Y)\rightarrow\operatorname{SO}^{+}(q,{\mathbb{Z}})$ is {±$\{\pm 1\}$ by Remark A.6.
@@
It remains to consider the middle entries of (B.2) and (B.3). We analyze when an element of $\beta(H^{\vee}\otimes H)\in{\mathbb{Z}}$ extends to map $f:\operatorname{H}^{2}(Y,{\mathbb{Z}})\rightarrow\operatorname{H}^{2}(Y,{\mathbb{Z}})$ which is ±$\pm\operatorname{id}$ on $T(Y)$. We assume first that it extends as $+\operatorname{id}$. Then $f$ has the form $x\mapsto x+\alpha\langle H,x\rangle H$ where $\alpha\in{\mathbb{Z}}$ because of the unimodularity of $\operatorname{H}^{2}(Y,{\mathbb{Z}})$.
Expressing $f(H)=\beta H$ yields $1+2\delta\alpha=\beta$, so $(2\delta)|(\beta-1)$. Applying this with $\beta=\frac{\delta}{e}bc+ead$ and recalling that $ead-\frac{\delta}{e}bc=1$ by the definition of an Atkin-Lehner element, we see that this implies $e|bc$, but this contradicts $ead-\frac{\delta}{e}bc=1$ unless $e=1$. Similarly, $f$ extends to $\operatorname{H}^{2}(Y,{\mathbb{Z}})$ as $-1$ on $T(Y)$ if and only if $e=\delta$. ∎

### v4

@@ 7.1. $Y\subset{\mathbb{P}}^{3}$ (degree 4)
$\otimes{\mathcal{O}}_{Y}(1)\in\operatorname{Aut}{\mathcal{D}}^{b}(Y).$
@@
In particular, $\Gamma_{0}^{+}(2):G]=4$ and it is easy to see that $G$ is the free product of $4$ copies of ${\mathbb{Z}}/2{\mathbb{Z}}$, generated by $s_{0},ts_{0}t^{-1},t^{2}s_{0}t^{-2}$ and $t^{3}s_{0}t^{-3}$. These elements correspond to the spherical twists with respect to exactly the restrictions of the vector bundles of the exceptional collection $\mathcal{F}$}. Figure 1 depicts a fundamental domain for $G$ which contains the corresponding $4$ fixed points of these generators.
@@ 7.2. $Y\subset Q_{3}$ (degree 6)
The group $\Gamma_{0}^{+}(3)\cong\operatorname{SO}^{+}(q,{\mathbb{Z}})$ has one elliptic point of order $2$, one elliptic point of order $6$ and one cusp ([16], Table 4). $X_{0}^{+}(3)$ has genus 0. Again, let $s_{0}\in\Gamma_{0}^{+}(3)$ be the transformation corresponding to $T_{{\mathcal{O}}_{Y}}\in\operatorname{Aut}{\mathcal{D}}^{b}(Y)$ and $t\in\Gamma_{0}^{+}(3)$ corresponding to $\otimes{\mathcal{O}}_{Y}(1)\in\operatorname{Aut}{\mathcal{D}}^{b}(Y)$. Then, similarly to the previous example, by Proposition 3.13 we have $(ts_{0})^{3}=s$, where $s$ denotes the transformation corresponding to $T_{{\mathcal{S}}^{*}|_{Y}}$. Hence $ts_{0}$ generates an elliptic subgroup of order $6$. Same as in the previous example, up to an even shift, the autoequivalence ${\mathcal{O}}_{Y}(1)\otimes T_{{\mathcal{O}}_{Y}}(-)\in\operatorname{Aut}{\mathcal{D}}^{b}(Y)$ has order $6$ as well. The subgroup $G$ generated by reflections has index $3$ in $\Gamma_{0}^{+}(3)$:
@@ 7.3. $Y\subset V_{5}$ (degree 10)
$1\to G\to\Gamma_{0}^{+}(5)\cong({\mathbb{Z}}/2{\mathbb{Z}})^{\ast 3}\xrightarrow{s_{0}\mapsto 0,s_{1}\mapsto 0,t^{2}\mapsto 0}{\mathbb{Z}}/2{\mathbb{Z}}\to 1.$
@@ 7.4. $Y\subset V_{22}$ (degree 22)
The group $\Gamma_{0}^{+}(11)\cong\operatorname{SO}^{+}(q,{\mathbb{Z}})$ has $4$ elliptic points of order $2$ and one cusp ([16], Table 4). $X_{0}^{+}(11)$ has genus 0.
@@ 7.5. Double cover of ${\mathbb{P}}^{2}$ (degree 2)
One observes that $\Gamma_{0}^{+}(1):G]=3$ and $G\cong{\mathbb{Z}}/2{\mathbb{Z}}\ast{\mathbb{Z}}/2{\mathbb{Z}}\ast{\mathbb{Z}}/2{\mathbb{Z}}$, generated by $s_{0}$, $ts_{0}t^{-1}$ and $t^{2}s_{0}t^{-2}$. These elements correspond to spherical twists with respect to the pullbacks of the bundles forming the standard exceptional collection on ${\mathbb{P}}^{2}$. Figure 5 shows a fundamental domain for $G$ containing the corresponding $3$ fixed points.
@@ 7.6. An example of an infinitely generated group of spherical twists (degree 8)
The group $\mathcal{G}$} can be finitely generated even if the K3 surface in question is an anticanonical section of a Fano threefold without a full exceptional collection. For instance, let $X$ be a section of the Plücker embedding of $\operatorname{Gr}(2,6)$ by a subspace of codimension $5$. Then $X$ is a Fano threefold of Picard rank $1$, index $1$ and degree $14$ and a smooth anticanonical section $Y\subset X$ is a K3 surface of degree $14$. For such $Y$ the group $\mathcal{G}$} is finitely generated by Corollary 5.11, because $\Gamma_{0}^{+}(7)$ has genus 0, 22 elliptic points of order $2$, which are both root points, $1$ elliptic point of order $3$ and $1$ cusp ([16, Table 4]). On the other hand, $X$ has no full exceptional collection since its Hodge diamond is not diagonal ($h^{2,1}=5$, see [26, §12.2]).
@@ 8. Exceptional collections on Fano threefolds
Let $X$ be a Fano threefold with Picard rank $1$ and very ample anticanonical class. Let $E,E^{\prime}$ be exceptional vector bundles on $X$ with the same restriction to a smooth anticanonical divisor $Y$. Assume that $\operatorname{Ext}^{1}(E,E\otimes\omega_{X}^{-1})=0$. Then $E\cong E^{\prime}$.
Remark 8.2.
Theorem 1.2 in the actual text [44] assumes that $E$ and $E^{\prime}$ have same class in $K_{0}(X)$ rather than the same restriction. However, examining Polishchuk’s proof one can see that the first step is to derive $E_{Y}=E^{\prime}_{Y}$ from this condition and afterwards the fact that $E$ and $E^{\prime}$ have the same class in $K_{0}(X)$ is not used.
@@
By Lemma 3.11 we have $i^{*}E\cong i^{*}F[n]$. By shifting $E$ we may assume without loss of generality assume that $n=0$. Having done this it remains to show that $E$ has one nonzero cohomology sheaf in degree 0. Consider the distinguished triangle
@@
    Since adbcad-bc is a unit and $a,d\in{\mathbb{Z}}_{p}$ we must have $v_{p}(b)+v_{p}(c)\geq 0$. Assume first $v_{p}(b)+v_{p}(c)>0$.
@@
    We now get that k=2m1k=2m-1 (in particular, $k\neq 0$ and hence $p\mid\delta$) and $v_{p}(c)=m$, $v_{p}(b)=1-m$. So
@@ Appendix B Understanding $\operatorname{Aut}^{+}\operatorname{H}^{\ast}(Y)$
[22, Corollary 15.2.12] A K3 surface $Y$ of Picard rank $1$ has a non-trivial automorphism group if and only if it has degree $2$. In the latter case there is a (unique) nontrivial automorphism $\iota\in\operatorname{Aut}(Y)$, which is the covering involution. The involution $\iota$ acts as the identity on ${\mathcal{N}}(Y)$ and as 1-1 on the transcendental lattice $T(Y)={\mathcal{N}}(Y)^{\perp}\subset H^{2}(Y,{\mathbb{Z}})$.
@@
Let $Y$ be a K3 surface with Picard rank 1 and $d\neq 2$. Then there is no element $\iota$ in $\operatorname{Aut}^{+}H^{\ast}(Y,{\mathbb{Z}})$ which acts as the identity on ${\mathcal{N}}(Y)$ and as 1-1 on $T(Y)$.
@@
Assume $\iota$ exists. It follows from the definition that $\iota$ preserves the grading on $H^{\ast}(Y,{\mathbb{Z}})$. In particular it acts on $H^{2}(Y,{\mathbb{Z}})$ such that $\iota(H)=H$ and $\iota|_{T(Y)}=-1$. It follows that $\iota$ is a reflection on $H^{2}(Y,{\mathbb{Z}})$ corresponding to $H$ and so it is given by the standard formula
@@
Let $Y$ be a K3 surface of odd Picard rank. The only Hodge isometries of the transcendental lattice $T(Y)$ are $\pm{\operatorname{id}}$.
@@
An element of $\operatorname{Aut}^{+}H^{*}(Y,{\mathbb{Z}})$ acts on the transcendental lattice $T(Y)$ by $\pm\operatorname{id}$, according to Proposition B.4. Since $\pm{\operatorname{id}}$ does not change the orientation of the positive 2-planes in $T(Y)_{\mathbb{R}}$, the map (B.1) is indeed well-defined.
If $\phi$ is a non-trivial element in the kernel of (B.1) then it is the identity on ${\mathcal{N}}(Y)$, and it acts as 1-1 on $T(Y)$. Invoking Proposition B.1 and Lemma B.3 finishes the proof. ∎
@@
which acts as $\pm 1$ on $T(Y)$. Note that by Lemma B.5 such an extension is unique for $d>2$ and for $d=2$ there is a ${\mathbb{Z}}/2{\mathbb{Z}}$ ambiguity which is inconsequential for what follows.
@@
    The kernel of $\operatorname{Aut}^{+}{\mathcal{N}}(Y)\rightarrow\operatorname{SO}^{+}(q,{\mathbb{Z}})$ is $\{\pm 1\}$ by Remark A.6.
@@
It remains to consider the middle entries of (B.2) and (B.3). We analyze when an element of $\beta(H^{\vee}\otimes H)\in{\mathbb{Z}}$ extends to map $f:\operatorname{H}^{2}(Y,{\mathbb{Z}})\rightarrow\operatorname{H}^{2}(Y,{\mathbb{Z}})$ which is $\pm\operatorname{id}$ on $T(Y)$. We assume first that it extends as $\operatorname{id}$. Then $f$ has the form $x\mapsto x+\alpha\langle H,x\rangle H$ where $\alpha\in{\mathbb{Z}}$ because of the unimodularity of $\operatorname{H}^{2}(Y,{\mathbb{Z}})$.
Expressing $f(H)=\beta H$ yields $1+2\delta\alpha=\beta$, so $(2\delta)|(\beta-1)$. Applying this with $\beta=\frac{\delta}{e}bc+ead$ and recalling that $ead-\frac{\delta}{e}bc=1$ by the definition of an Atkin-Lehner element, we see that this implies $e|bc$, but this contradicts $ead-\frac{\delta}{e}bc=1$ unless $e=1$. Similarly, $f$ extends to $\operatorname{H}^{2}(Y,{\mathbb{Z}})$ as 1-1 on $T(Y)$ if and only if $e=\delta$. ∎